Operations Report – February 20th

Crew 189 Operations Report 20Feb2018
SOL: 2

Name of person filing report: PAYEN Gabriel

Non-nominal systems: Backpack 3


Solar panels < 70% without generator running

Notes on non-nominal systems:

Backpack 3 :

All engineering checks were nominals, but during the EVA both fans turned out of service. Also, the necklace which holds the helmet was not stretched enough so that the helmet fell on the neck of Gabriel during half of the EVA.

Backpack 3 joined backpack 4 to the list of backpacks the crew need to fix. We’ll certainly begin to fix them tomorrow.

WiFi :

Only 15 Mo were left this morning (immediately after the very connection of the day were established). Yet, the leak is unknow. I reported the problem this morning and the crew have communicated with CapCom to find a solution.

Solar panels :

I had to go out of the SIM to turn on the generator at 6:20 PM. Solar SOC was at 53%, I hope batteries are not injured.

Generator (hours run): Turned of at 12:00 and turned on at 6:20PM.

Solar— SOC % (Before generator is run at night) /!\ 53%

Diesel – 70%

Propane – 40 psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 7 gallons.

Water (trailer) – 0 gallons.

Water (static) – 65% gallons

Trailer to Static Pump used – NO

Water (loft) – Static to Loft Pump used – YES

Water Meter: 0.26m

Toilet tank emptied: NO

ATV’s Used: NO

Deimos rover used: YES

Hours: 109.1

Beginning charge: 100%

Ending charge: 100% (the EVA team doesn’t travel a lot)

Currently charging: YES

Spirit rover used: NO

Hours: 20.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: NO

Opportunity rover used: YES

Hours: 16.6

Beginning charge: 100%

Ending charge: 45%

Currently charging: YES

Curiosity rover used: NO

Hours: 100%

Beginning charge: –

Ending charge: –

Currently charging: NO

General notes and comments: –

Summary of internet:

-Still seeking the leak.

Summary of suits and radios:

-Backpack 3 is put aside. Starting fixing backpack 3 and 4 tomorrow.

Summary of Hab operations:

-Two thermometers and the shape for the GreenHab received in the airlock. Jeremy (GreenHab officer) will set it up tomorrow.

-The pipe in the low desk seems not to flow anymore. All water devices in the Hab and GreenHab were tested one by one and none trigger the leak.

-Considering a Dr Rupert’s request, a mattress will be put in the airlock between 8PM and 9 PM.

Summary of GreenHab operations: –

Summary of ScienceDome operations: –

Summary of RAMM operations: –

Summary of health and safety issues: –

Questions, concerns and requests to Mission Support: –

Copyright © The Mars Society. All rights reserved. | Main Site